OpenStudy (anonymous):

The cost for admission to the zoo is twice as much for adults as for children. The total price of admission to the zoo for two adults and eight children is $48. Find the price of admission to the zoo for a child.

OpenStudy (anonymous):

if you call x= cost of admission for child, adult = 2x right?

OpenStudy (anonymous):

so, 2 adults = 2*(2x) = 4x 8 children = 8x add and =48 solve it for x = your answer

OpenStudy (anonymous):

did you get the answer?

OpenStudy (anonymous):

is that it? could you set it up?

OpenStudy (anonymous):

4x+8x=48 ...try now :)

OpenStudy (anonymous):

ok what do i do? or do i just solve for x?

OpenStudy (anonymous):

12x=48 x=...

OpenStudy (anonymous):

remember, that x= price of child admission - exactly what you need!

OpenStudy (anonymous):

o ok sweet thanks br0!

OpenStudy (anonymous):

x=4 Welcome!!